Use the four-step process to find the slope of the tangent line to the graph of the given function at any point. (Simplify your answers completely.)
f(x) = 9 − 7x
Step 1: f(x + h)=
Step 2: f(x + h) − f(x)=
Step 3: f(x + h) − f(x)/h=
Step 4: f '(x) = lim h→0 f(x + h) − f(x)/h=

Answer by Boreal(15235)

f(x+h)=9-7x-7h
f(x+h)-f(x)=9-7x-7h-9+7x
=-7h
divide that by h=-7
f'(x)=lim h goes to 0 of- 7=-7